# StableMaterials: Enhancing Diversity in Material Generation via Semi-Supervised Learning

### v3

@@ 3.4. Fast High-Resolution Generation
To improve generation speed, we fine-tune a Latent Consistency Model (LCM) [Luo et al., 2023]. LCM performs a one-stage guided distillation of an augmented Probability Flow ODE (PF-ODE) and directly predicts the solution at $t=0$ through a consistency function $f_{\theta}(z_{t},c,t)\mapsto z_{0}$. Unlike two-stage methods [Meng et al., 2023], LCM integrates the guidance scale $\omega$ directly into the PF-ODE and uses a skip-step strategy, ensuring consistency between time steps $t_{n+k}$ and $t_{n}$. This design avoids alignment issues present in typical two-stage approaches. As a result, LCM enables generation in only a few steps, resulting in faster material synthesis.
@@
Noise rolling [Vecchio et al., 2024a] has been used to achieve tileability through iterative diffusion but becomes less effective with fewer steps. We address this limitation proposing the features rolling, which shifts rolling from the noisy inputs to the U-Net features. Within each convolution and attention layer, we randomly roll and then reverse the feature maps, thus preserving edge continuity while requiring fewer diffusion steps. For highly structured materials, we enable features rolling only after the first diffusion step to retain their global layout. We compare features rolling with other tileability methods in the Supplemental Materials.
@@
Combining features rolling with patch-based diffusion provides efficient, high-resolution synthesis, but patch-wise generation alone can introduce inconsistencies across tiles. We address this with a two-stage pipeline, similar to SDXL [Podell et al., 2023], using SDEdit [Meng et al., 2021] for refinement. Specifically, we first generate at $512\times 512$ resolution, then refine the output to the desired resolution with a strength of $0.5$, balancing new high-frequency details against global consistency. The refiner model is trained on full-resolution $512\times 512$ crops from 4K materials (no downsampling), ensuring it captures fine surface details. We demonstrate the effectiveness of this approach in Sec. 4.4.3, with additional samples provided in the Supplemental Materials.
4. Implementation & Results
In this section, we first introduce the datasets used in our work, we then assess the generation capabilities of StableMaterials and compare it with recent state-of-the-art methods. Finally, we evaluate our design choices in the ablation studies.
4.1. Dataset
We train our model on the combined MatSynth [Vecchio and Deschaintre, 2024] and Deschaintre et al. [2018] datasets, for a total of $6,198$ unique PBR materials, using the original training/test splits. Our dataset includes 5 material maps (Basecolor, Normal, Height, Roughness, Metallic) and their renderings under different environmental illuminations. We complement the dataset with $4,000$ texture-text pairs from SDXL [Podell et al., 2023], using $200$ prompts. We query a ChatGPT [OpenAI, 2024] model to suggest relevant material prompts. The full list of prompts used, as well as samples generated from SDXL, are provided in the Supplemental Material.
4.2. Technical details
@@
Autoencoder
The compression model is trained with mini-batch gradient descent, using the Adam optimizer [Kingma and Ba, 2014] and a batch size of $8$. We train the model for $1,000,000$ iterations with a learning rate of $4.5\cdot 10^{-4}$ and enable the $\mathcal{L}_{\text{adv}}$ after $300,000$ iterations as in Esser et al. [2021]. We first train a multi-encoder VAE [Vecchio et al., 2024b]; then we fine-tune a single-encoder model to compress the concatenated maps into the same disentangled latent space, reducing the total of parameters from 271M to 101M, while keeping similar reconstruction performance (Tab. 2). We fine-tune both the single encoder model and the texture encoder for $100,000$ steps while keeping the decoder frozen.
Latent Diffusion model
The diffusion model is trained supervisedly for $400,000$ iterations with a batch size of $16$ using an AdamW [Loshchilov and Hutter, 2017] optimizer, with a learning rate of $3.2\cdot 10^{-5}$. The model is fine-tuned semi-supervisedly for $200,000$ iterations with a batch size of $8$. We fine-tune the refiner model for $50,000$ iterations using a batch size of $16$ and a learning rate of $2\cdot 10^{-7}$. Both models use the original OpenAI U-Net architecture with 18 input and output channels.
Latent Consistency model
The latent consistency model [Luo et al., 2023] is fine-tuned for $10,000$ iterations, with a batch size of $16$, using an AdamW optimizer, with a learning rate value of $1\cdot 10^{-6}$. We use a linear schedule for $\beta$ and denoise using the LCM (Latent Consistency Models) sampling schedule at inference time with $T=4$ steps.
Inference
@@
Figure 4. Image-prompting. StableMaterials accurately captures the visual appearance of the input image, producing realistic materials for both in-domain (on the left) and out-domain (on the right) prompts. The render highlights the model’s ability to handle diverse and complex surfaces.
@@
Figure 5. Text-prompting. StableMaterials closely follow the input prompt, producing realistic materials for both in-domain (on the left) and out-domain (on the right) samples. The render highlights the model’s ability to generate accurate properties for different types of materials.
@@
Figure 6. Comparison for image-prompting. We compare StableMaterials with MatFuse, MatGen, and MaterialPalette in image-prompted generation, showing two in-domain (left column) and two out-domain (right column) renderings per model. StableMaterials improves over previous methods quality and ability to captures the visual appearance of the input image. Full set of maps in Supplemental Materials.
@@
Figure 7. Comparison for text prompting. We compare StableMaterials with MatFuse, MatGen, and Substance 3D Sampler on text-prompted generation, showing two in-domain (left column) and two out-domain (right column) renderings per model. StableMaterials better follows the input prompt and successfully models out-domain materials. Visual quality is on par or better than models trained on larger datasets. Full set of maps in Supplemental Materials.
4.3. Results and comparison
All results show both generated material maps and ambient-lit renderings. For evaluation, we carefully selected test prompts to include both in-domain concepts (present in training categories) and novel out-domain concepts, ensuring no direct overlap with training prompts. Additional samples, material editing results, and a CLIP-based nearest-neighbor search in our training database are included in the Supplemental Materials.
@@ 4.3. Results and comparison
We compare StableMaterials against MatFuse [Vecchio et al., 2024b], MatGen [Vecchio et al., 2024a], Material Palette [Lopes et al., 2024], and Adobe Substance 3D Sampler [Adobe, 2024], as shown in Figures 6 and 7 for image and text prompting, respectively. MatFuse, Material Palette, and StableMaterials each use public datasets, while MatGen and Sampler rely on private data. MatFuse is limited to a $256\times 256$ resolution, often generating blurry or simplistic outputs and struggling with complex textures. Material Palette and Sampler take a two-step approach (texture generation followed by SVBRDF estimation).This approach benefits from leveraging large image models, but suffers from biases inherent in SVBRDF predictions, which guess material properties from light-surface interactions, that are further exacerbated by the nonmaterial-specific training sometimes generating natural images (with perspective) instead of surfaces. Moreover, MaterialPalette fine-tunes a LoRA for each prompt, introducing a time and computation overhead. MatGen, on the other hand, produces high-quality materials, but presents over-sharpening artifacts and struggles in following more complex promps. In contrast, StableMaterials directly outputs PBR maps using a tailored latent representation and a two-stage generation process, mitigating resolution constraints and reducing artifacts. We only show the generation renderings in Figures 6 and 7, with the PBR maps and additional samples included in the Supplemental Materials.
@@
Material quality is difficult to assess with standard image metrics (FID [Heusel et al., 2017], IS [Salimans et al., 2016]) due to the different data distribution from that of natural images. Instead, we leverage CLIP-based metrics (CLIP Score [Hessel et al., 2021] and CLIP-IQA [Wang et al., 2023]), which assess both semantic alignment with prompts and perceptual quality of outputs. Table 1 shows StableMaterials either outperforms or matches state-of-the-art methods—some trained on larger datasets—and significantly improves over MatFuse. Comparison is carried out on 80 text-conditioned generation.
Table 1. Quantitative comparison. We compare the generation quality of StableMaterials to MatFuse, MatGen and Substance 3D Sampler. The CLIP-IQA score is computed using the “high-quality/low-quality” contrastive pair.
@@
4.4. Ablation Study
@@
4.4.1. VAE Architecture
@@
Basecolor Normal Height Rough. Metallic
$-\mathcal{E}$ (271M par.) 0.030 0.035 0.030 0.032 0.016
$-\mathcal{E}$ (101M par.) 0.028 0.037 0.030 0.032 0.015
@@ 4.4.2. Training strategies
Figure 8. Ablation study of the training stategies. The model effectively generates high-quality material with image or text prompts but struggles with unrepresented materials. Semi-supervised learning improves generation quality and diversity, including new materials.
@@
Figure 9. Ablation studies comparing (a) the effect of refinement on quality and (b) different approaches to achieving high-resolution. Results show that the use of a diffusion refiner significantly enhances generation quality and sharpness, and two-stage approach generates at the model’s native resolution avoids the inconsistency in scale of the patched diffusion.
4.4.3. High-Resolution Generation
5. Limitations and Future Work
@@
Figure 10. Limitations. Left to right: (1) Struggles with complex prompts describing spatial relations. (2) Unable to represent complex figures or patterns. (3) Can hallucinate reflectance properties.
6. Conclusion
We introduce StableMaterials, a novel diffusion-based model for fast, tileable, high-resolution material generation. By integrating semi-supervised training and knowledge distillation from large-scale pretrained models, StableMaterials overcomes the lack of annotated data and delivers greater realism and variety in PBR materials. Moreover, feature-rolling enables tileable generation in few-steps settings, making our approach practical for real-world applications.

### v4

@@ 3.4 Fast High-Resolution Generation
To improve generation speed, we fine-tune a Latent Consistency Model (LCM) [36]. LCM performs a one-stage guided distillation of an augmented Probability Flow ODE (PF-ODE) and directly predicts the solution at $t=0$ through a consistency function $f_{\theta}(z_{t},c,t)\mapsto z_{0}$. Unlike two-stage methods [41], LCM integrates the guidance scale $\omega$ directly into the PF-ODE and uses a skip-step strategy, ensuring consistency between time steps $t_{n+k}$ and $t_{n}$. This design avoids alignment issues present in typical two-stage approaches. As a result, LCM enables generation in only a few steps, resulting in faster material synthesis.
@@
Noise rolling [60] has been used to achieve tileability through iterative diffusion but becomes less effective with fewer steps. We address this limitation proposing the features rolling, which shifts rolling from the noisy inputs to the U-Net features. Within each convolution and attention layer, we randomly roll and then reverse the feature maps, thus preserving edge continuity while requiring fewer diffusion steps. For highly structured materials, we enable features rolling only after the first diffusion step to retain their global layout. We compare features rolling with other tileability methods in Supplemental Materials.
@@
Combining features rolling with patch-based diffusion provides efficient, high-resolution synthesis, but patch-wise generation alone can introduce inconsistencies across tiles. We address this with a two-stage pipeline, similar to SDXL [44], using SDEdit [40] for refinement. Specifically, we first generate at $512\times 512$ resolution, then refine the output to the desired resolution with a strength of $0.5$, balancing new high-frequency details against global consistency. The refiner model is trained on full-resolution $512\times 512$ crops from 4K materials (no downsampling), ensuring it captures fine surface details. We demonstrate the effectiveness of this approach in Sec. 4.4, with additional samples provided in Supplemental Materials.
4 Implementation & Results
In this section, we first introduce the datasets used in our work, we then assess the generation capabilities of StableMaterials and compare it with recent state-of-the-art methods. We evaluate our design choices in the ablation studies.
4.1 Dataset
We train our model on the combined MatSynth [58] and Deschaintre et al. [6] datasets, for a total of $6,198$ unique PBR materials, using the original training/test splits. Our dataset includes 5 material maps (Basecolor, Normal, Height, Roughness, Metallic) and their renderings under different environmental illuminations. We complement the dataset with $4,000$ texture-text pairs from SDXL [44], using $200$ prompts. We query a ChatGPT [43] model to suggest relevant material prompts. The full list of prompts and the SDXL textures, are provided in Supplemental Material.
4.2 Technical details
@@
Autoencoder is trained with mini-batch gradient descent, using the Adam optimizer [29] and a batch size of $8$. We train the model for $1,000,000$ iterations with a learning rate of $4.5\cdot 10^{-4}$ and enable the $\mathcal{L}_{\text{adv}}$ after $300,000$ iterations as in Esser et al. [10]. We first train a multi-encoder VAE [61]; then we fine-tune a single-encoder model to compress the concatenated maps into the same disentangled latent space, reducing the total of parameters from 271M to 101M, while keeping similar reconstruction performance (Tab. 2). We fine-tune both the single encoder model and the texture encoder for $100,000$ steps while keeping the decoder frozen.
@@
Figure 3: Image-prompting. StableMaterials accurately captures the visual appearance of the input image, producing realistic materials for both in-domain (on the left) and out-domain (on the right) prompts. The render highlights the model’s ability to handle diverse and complex surfaces.
@@
Figure 4: Text-prompting. StableMaterials closely follow the input prompt, producing realistic materials for both in-domain (on the left) and out-domain (on the right) samples. The render highlights the model’s ability to generate accurate properties for different types of materials.
@@
Figure 5: Comparison for image-prompting. We compare StableMaterials with MatFuse, MatGen, and MaterialPalette in image-prompted generation, showing two in-domain (left column) and two out-domain (right column) renderings per model. StableMaterials improves over previous methods quality and ability to captures the visual appearance of the input image.
@@
Figure 6: Comparison for text prompting. We compare StableMaterials with MatFuse, MatGen, and Substance 3D Sampler on text-prompted generation, showing two in-domain (left column) and two out-domain (right column) renderings per model. StableMaterials better follows the input prompt and successfully models out-domain materials. Visual quality is on par or better than models trained on larger datasets.
Latent Diffusion model is trained supervisedly for $400,000$ iterations with a batch size of $16$ using an AdamW [35] optimizer, with a learning rate of $3.2\cdot 10^{-5}$. The model is fine-tuned semi-supervisedly for $200,000$ iterations with a batch size of $8$. We fine-tune the refiner model for $50,000$ iterations using a batch size of $16$ and a learning rate of $2\cdot 10^{-7}$. Both models use the original OpenAI U-Net architecture with 18 input and output channels.
Latent Consistency model is fine-tuned for $10,000$ iterations, with a batch size of $16$, using an AdamW optimizer, with a learning rate value of $1\cdot 10^{-6}$. We use a linear schedule for $\beta$ and denoise using the LCM sampling schedule at inference time with $T=4$ steps.
Inference
4.3 Results and comparison
All results show both generated material maps and ambient-lit renderings. For evaluation, we carefully selected test prompts to include both in-domain concepts (present in training categories) and novel out-domain concepts, ensuring no direct overlap with training prompts. Additional samples, material editing results, and a CLIP-based nearest-neighbor search are included in the Supplemental Materials.
@@ 4.3 Results and comparison
We compare StableMaterials against MatFuse [61], MatGen [60], Material Palette [34], and Adobe Substance 3D Sampler [1], as shown in Figures 5 and 6 for image and text prompting, respectively. MatFuse, Material Palette, and StableMaterials each use public datasets, while MatGen and Sampler rely on private data. MatFuse is limited to a $256\times 256$ resolution, often generating blurry or simplistic outputs and struggling with complex textures. Material Palette and Sampler take a two-step approach (texture generation followed by SVBRDF estimation).This approach benefits from leveraging large image models, but suffers from biases inherent in SVBRDF predictions, which guess material properties from light-surface interactions, that are further exacerbated by the nonmaterial-specific training sometimes generating natural images (with perspective) instead of surfaces. Moreover, MaterialPalette fine-tunes a LoRA for each prompt, introducing a time and computation overhead. MatGen, on the other hand, produces high-quality materials, but presents over-sharpening artifacts and struggles in following more complex promps. In contrast, StableMaterials directly outputs PBR maps using a tailored latent representation and a two-stage generation process, mitigating resolution constraints and reducing artifacts. We only show the generation renderings in Figures 5 and 6, with the PBR maps and additional samples included in the Supplemental Materials.
User Study
To validate the quality of our results, we conducted a user study with 20 domain experts. Participants compared material generations from the different methods across ten prompts. For each prompt, they selected their preferred sample and rated all four materials on a 1–5 scale for prompt fidelity, realism, and material consistency. StableMaterials was preferred 105 times, compared to 66 for Sampler, 25 for MatGen, and 4 for MatFuse. It also achieved the highest average rating (3.50), outperforming Sampler (2.71), MatGen (2.28), and MatFuse (1.85). A one-way ANOVA confirmed the differences are statistically significant ($F=60.5$, $p<0.05$). A chi-square goodness-of-fit test on the preference counts $[4,25,66,105]$ against a uniform distribution yielded $\chi^{2}=120.44$, $p<0.001$, indicating strong deviation from chance. A correlation coefficient of $R=0.99$ confirms alignment between average ratings and preference counts.
@@
Material quality is difficult to assess with standard image metrics (FID [20], IS [50]) due to the different data distribution from that of natural images. Instead, we leverage CLIP-based metrics (CLIP Score [19] and CLIP-IQA [63]), which assess both semantic alignment with prompts and perceptual quality of outputs. Table 1 shows StableMaterials either outperforms or matches state-of-the-art methods—some trained on larger datasets—and significantly improves over MatFuse. Comparison is carried out on 80 text-conditioned generation.
@@
Table 1: Quantitative comparison. We compare the generation quality of StableMaterials to MatFuse, MatGen and Substance 3D Sampler. The CLIP-IQA score is computed using the “high-quality/low-quality” contrastive pair.
4.4 Ablation Study
@@
VAE Architecture
Albedo Normal Height Rough. Metal.
Multi-$\mathcal{E}$ (271M par.) 0.030 0.035 0.030 0.032 0.016
Single-$\mathcal{E}$ (101M par.) 0.028 0.037 0.030 0.032 0.015
@@
High-Resolution Generation
5 Limitations and Future Work
@@
Figure 7: Ablation study of the training stategies. The model effectively generates high-quality material with image or text prompts but struggles with unrepresented materials. Semi-supervised learning improves generation quality and diversity, including new materials.
@@
Figure 8: Ablation studies comparing (a) the effect of refinement on quality and (b) different approaches to achieving high-resolution. Results show that the use of a diffusion refiner significantly enhances generation quality and sharpness, and two-stage approach generates at the model’s native resolution avoids the inconsistency in scale of the patched diffusion.
@@
Figure 9: Limitations. Left to right: (1) Struggles with complex prompts describing spatial relations. (2) Unable to represent complex figures or patterns. (3) Can hallucinate reflectance properties.
6 Conclusion
We introduce StableMaterials, a novel diffusion-based model for fast, tileable, high-resolution material generation. By integrating semi-supervised training and knowledge distillation from large-scale pretrained models, it overcomes the lack of annotated data and delivers greater realism and variety in PBR materials. Moreover, features rolling enables tileable generation in few-steps settings, making our approach practical for real-world applications.